# List strong and list normal edge-coloring of (sub)cubic graphs

### v2

@@ Abstract
It turns out that similarly to the strong edge-coloring, the list normal edge-coloring is much more restrictive and consequently for many graphs the list normal chromatic index is greater than the normal chromatic index. In particular, we show that there are cubic graphs with the list normal chromatic index at least $9$, there are bridgeless cubic graphs with its value at least $8$, and there are cyclically $4$-edge-connected cubic graphs with value at least $7$.
@@ 1.1 List strong edge-coloring
The second question, we consider in regard to the list strong edge-coloring, is whether the values of the list strong chromatic index and the strong chromatic index of subcubic graphs coincide. In particular, we are interested in a special case of the question proposed by Dai et al. [5].
@@
The motivation for the question comes from the List (Edge) Coloring Conjecture stating that the values of the chromatic index and the list chromatic index of any graph coincide. The conjecture was stated independently by several researchers (see [20, Problem 12.20] for more details) and in general it is still widely open; cf., e.g., [3] for a short survey.
@@
Interestingly, there are also some planar graphs (e.g., the dodecahedron) and bipartite graphs (e.g., the generalized Petersen graph $GP(10,3)$) among the graphs with different values for the two invariants. Finally, for the case of the Petersen graph, we prove the exact value of the list strong chromatic index.
@@ 1.2 List normal edge-coloring
In this paper, in relation to the list strong edge-colorings, we also study the properties of the list version of the normal edge-coloring. For a cubic graph $G$, the list normal edge-coloring and the list normal chromatic index, $\chi_{n,l}^{\prime}(G)$, are defined analogously to the list strong variants.
@@
We show that, similarly as for the list strong edge-coloring, also in the list normal edge-coloring there are graphs $G$ with $\chi_{n,l}^{\prime}(G)>\chi_{n}^{\prime}(G)$. In particular, there is an infinite family of cubic graphs with the list normal chromatic index at least $9$, there are bridgeless cubic graphs with the list normal chromatic index at least 8, and there is an infinite family of cyclically 4-edge-connected cubic graphs with the list normal chromatic index at least 7. Interestingly, our examples of bridgeless graphs for the above results are all from class I, and therefore they all have the normal chromatic index equal to $3$.
@@ 3 Upper bound on the list strong chromatic index
By the minimality of $G$, there exists a list strong edge-coloring $\varphi^{\prime}$ of $G^{\prime}=G\setminus\left\{v_{1},v_{2},v_{3},v_{4},v_{5}\right\}$ for any list assignment $L$ with lists of size at least $10$. Let $\varphi$ be the coloring of $G$ induced by $\varphi^{\prime}$. Then, only the edges of $C$ and the pendant edges except $y_{0}$ are non-colored in $\varphi$. The edges $x_{0}$ and $x_{5}$ have at least $5$ available colors, the edges $x_{1},x_{2},x_{3},x_{4}$ have at least $6$, $y_{1}$ and $y_{5}$ have at least $3$, and $y_{2},y_{3},y_{4}$ have at least $4$ available colors.
@@
$P_{C_{k}}^{0}(x_{0},x_{1},y_{0},y_{1})=\mathrm{coef}(P_{C_{k}}^{k-1};\ x_{k-1}^{4}y_{k-1}^{3})\cdot(x_{0}-x_{1})(x_{0}-y_{0})(x_{0}-y_{1})(x_{1}-y_{0})(x_{1}-y_{1})(y_{0}-y_{1})\,,$
@@ 4 Graphs $G$ with $\chi_{s}^{\prime}(G)<\chi_{s,l}^{\prime}(G)$
In this section, we give a negative answer to Question 1.5 by proving Theorem 1.6. First, we recall the result about cubic graphs with the strong chromatic index equal to $5$. It uses a notion of covering graph defined as follows. A surjective graph homomorphism $f\colon\tilde{G}\to G$ is called a covering projection if for every vertex $\tilde{v}$ of $\tilde{G}$ the set of edges incident with $\tilde{v}$ is bijectively mapped onto the set of edges incident with $f(\tilde{v})$. The graph $G$ is usually referred to as the base graph and $\tilde{G}$ as a covering graph or a lift of $G$. A graph $\tilde{G}$ covers $G$ if there exists such a covering projection.
@@
Let $G^{\prime}$ be the graph obtained from $G$ be removing all the edges of $P^{\prime}$ except $uv_{1}$ and $vv_{2}$. Clearly, $G^{\prime}$ is the graph $R$ with one edge removed and replaced with two pendant edges, and thus it admits a strong $5$-edge-coloring $\varphi^{*}$ induced by the coloring $\varphi$ of $R$. Note that in $\varphi$, the edges $uv_{1}$ and $vv_{2}$ receive the same color (the color $\varphi(uv)$). Now, we show that in any strong $5$-edge-coloring of $G^{\prime}$ these two edges must be colored with the same color. Since the edges of $G^{\prime}$ in $L$ have the same lists of size $5$, this will imply that the two edges must receive the same color in any strong $L$-edge-coloring.
First, observe that in $\varphi^{*}$ the only common color the vertices $u$ and $v$ are incident with is color $a=\varphi(uv)$. Let $b\neq a$ be a color incident with $u$, and $c\neq a$ a color incident with $v$. Let $k$ be the number of edges of color $b$ in $\varphi^{*}$. Since every edge of color $b$ is adjacent to edges of all other four colors, the edges of $G^{\prime}$ colored with $b$ in $\varphi^{*}$ together with their adjacent edges cover all the edges of $G^{\prime}$ (every edge exactly once) except $vv_{2}$; we denote this (almost) covering $C_{b}$. Similarly, we the edges of $G^{\prime}$ colored with $c$ in $\varphi^{*}$ together with their adjacent edges cover all the edges of $G^{\prime}$ (every edge exactly once) except $uv_{1}$; we denote this (almost) covering $C_{c}$.
@@
Similarly, we can show that around the vertices $u$ and $v$ in $G$, all five colors appear (i.e., the only common incident color is the color of the edges $uv_{1}$ and $vv_{2}$). Observe that the edges of $G^{\prime}$ colored with $a$ in $\varphi^{*}$ (except the edges $uv_{1}$ and $vv_{2}$) together with their adjacent edges cover all the edges of $G^{\prime}$ (every edge exactly once) except the edges incident with $u$ and $v$; we denote this covering $C_{a}$. Again, in every strong $5$-edge-coloring of $G^{\prime}$, on the edges of $C_{a}$ every color appears $k-1$ times, while in the whole graph every color appears on $k$ edges, except the color of $uv_{1}$ and $vv_{2}$, which appears $k+1$ times. This means, that $u$ and $v$ together see all five colors.
Now consider the coloring of the edges of $P^{\prime}$. Clearly, in any strong $L$-edge-coloring, all the five colors from $\left\{1,2,3,4,5\right\}$ appear on the inner edges of $P^{\prime}$. Similarly, since every spoke edge of $P^{\prime}$ sees $4$ distinct colors on the inner edges of $P^{\prime}$, every spoke edge can be colored with precisely one of the colors from $\left\{1,2,3,4\right\}$ or color $6$, except for the spoke edge that does not have color $5$ in its $2$-edge-neighborhood—that edge must be colored with $6$.
There are three non-isomorphic possibilities on which inner edge color $5$ appears. First, suppose that $u_{1}u_{4}$ is colored with $5$. Then, it is easy to see that $u_{5}v_{5}$ and $v_{2}v_{3}$ must be colored with $6$, meaning that no outer edge of $P^{\prime}$ can be colored with $5$, and we cannot color all edges of $G$.
Second, suppose that $u_{3}u_{5}$ is colored with $5$. Similarly as above, we can deduce that $u_{4}v_{4}$, $uv_{1}$, and $vv_{2}$ must be colored with $6$, meaning that we cannot color any outer edge with $5$, and thus we cannot color all edges of $G$.
So, we may assume that $u_{1}u_{3}$ is colored with $5$. Then, we can deduce that $u_{2}v_{2}$ and $v_{1}v_{5}$ must be colored with $6$. This means that, since $uv_{1}$ and $vv_{2}$ have the same color, some color, different from the color of $uv_{1}$, must be incident with $u$ and $v$. As we showed above, this is not possible, and therefore a strong $L$-edge-coloring of $G$ does not exist. ∎
@@
Now, we consider several cases regarding the possible colorings of the maximum induced matchings. Note that throughout the process of coloring the edges, as soon as some color is picked for an edge $e,\ e\in M_{i}$, this color is removed from the lists of the edges which are in conflict with $e$, i.e., the edges of the maximum induced matchings different from $M_{i}$.
@@ 5 Graphs $G$ with $\chi_{n}^{\prime}(G)<\chi_{n,l}^{\prime}(G)$
In this section, we consider the results on list normal edge-coloring. As already mentioned, lists of size at least $10$ are always enough to find a normal list edge-coloring of a cubic graph. We do not know whether this bound is tight; currently, there are only examples of graphs with the list normal chromatic index equal to $9$.
@@
There is an infinite family of cubic graphs with the list normal chromatic index at least $9$.
@@
There are bridgeless cubic graphs with the list normal chromatic index at least $8$.
@@
The graph in the proof of Theorem 5.2 has a $2$-edge-cut. So, again a question arises whether there are cubic graphs with high list normal chromatic index and high connectivity. We focused on cyclically $4$-edge-connected cubic graphs and surprisingly there are such graphs with the list normal chromatic index at least $7$.
@@
There is an infinite family of cyclically $4$-edge-connected cubic graphs with the list normal chromatic index at least $7$.
@@ 6 Conclusion
One of the main results of this paper is the tight upper bound of $10$ colors for the strong chromatic index of subcubic graphs. However, since this bound is only known to be attained by the Wagner graph and graph containing the $K_{3,3}$ with a subdivided edge as a subgraph, one may ask whether there are other examples of such graph, perhaps with smaller chromatic index.
@@
As the second main result, we proved that the strong chromatic index and the list strong chromatic index differ for some graphs; we provided an infinite family of such graphs, but the family only contains graphs with the minimum possible value of the strong chromatic index, and it does not seem likely that for graphs with the strong chromatic index closer to the general upper bound of $10$ colors, their list strong chromatic index will be different. Therefore, we propose a rather bold statement, which is in line with Conjecture 1.3.
@@
The first step towards proving this conjecture would be proving that the list strong chromatic index of any connected bridgeless subcubic graph, not isomorphic to the Wagner graph, is at most $9$. Or even more specifically, finding the exact upper bounds for the list strong chromatic indices of special graph families such as planar graphs and bipartite graphs would also give a relevant insight in the topic.
@@
In the case of the list normal edge-coloring, we have the upper bound given by Theorem 1.4, but we do not have an example of a graph attaining the bound; in fact, we do not believe one exists.
@@
Conjecture 1.8 assumes only bridgeless cubic graphs. We showed in Theorem 5.2 that in the list version, there are bridgeless cubic graphs with the list normal chromatic index at least $8$. In the proof of the theorem, we only provided one graph of order $12$. However, we are only aware of two other graphs with the list normal chromatic index at least $8$; namely the Wagner graph and the graph obtained from $K_{3,3}$ in which one vertex is truncated (see Figure 10). We also remark here without a proof that with some additional effort, one can show that the list normal chromatic index of the Wagner graph is equal to $8$.
Figure 10: Two bridgeless cubic graphs with the list strong chromatic index at least $8$. If the bold edges receive unique lists of seven colors and all the other edges the same lists of seven colors, then one can not realize a list normal edge-coloring.
Based our results and additional computer tests on small graphs, we confidently propose also the following.
@@
    $k$-edge-connected, for some integer $k$?

### v3

@@ Abstract
It turns out that similarly to strong edge-coloring, list normal edge-coloring is much more restrictive and consequently for many graphs the list normal chromatic index is greater than the normal chromatic index. In particular, we show that there are cubic graphs with list normal chromatic index at least $9$, there are bridgeless cubic graphs with its value at least $8$, and there are cyclically $4$-edge-connected cubic graphs with value at least $7$.
@@ 1.1 List strong edge-coloring
The second question regarding the list strong edge-coloring is whether the values of the list strong chromatic index and the strong chromatic index of subcubic graphs coincide. In particular, we are interested in a special case of the question proposed by Dai et al. [5].
@@
The motivation for the question comes from the List (Edge) Coloring Conjecture stating that the values of the chromatic index and the list chromatic index of any graph coincide. The conjecture was stated independently by several researchers (see [21, Problem 12.20] for more details) and in general it is still widely open; cf., e.g., [3] for a short survey.
@@
Interestingly, there are also some planar graphs (e.g., the dodecahedron) and bipartite graphs (e.g., the generalized Petersen graph $GP(10,3)$) among the graphs with different values for the two invariants. Note that the above results are independently obtained also by Hasanvand [13].
Finally, for the case of the Petersen graph, we prove the exact value of the list strong chromatic index.
@@
After publishing the preprint of this paper, we were notified by M. Hasanvand that the result of Theorem 1.7 follows also from the result of Kierstead on complete multipartite graphs [23].
@@ 1.2 List normal edge-coloring
In this paper, in relation to the list strong edge-colorings, we also study the properties of the list version of the normal edge-coloring. For a cubic graph $G$, list normal edge-coloring and the list normal chromatic index, $\chi_{n,l}^{\prime}(G)$, are defined analogously to the list strong variants.
@@
We show that, similarly to the list strong edge-coloring, also in the list normal edge-coloring there are graphs $G$ with $\chi_{n,l}^{\prime}(G)>\chi_{n}^{\prime}(G)$. In particular, there is an infinite family of cubic graphs with list normal chromatic index at least $9$, there are bridgeless cubic graphs with list normal chromatic index at least 8, and there is an infinite family of cyclically 4-edge-connected cubic graphs with list normal chromatic index at least 7. Interestingly, our examples of bridgeless graphs for the above results are all from class I, and therefore they all have the normal chromatic index equal to $3$.
@@ 3 Upper bound on the list strong chromatic index
By the minimality of $G$, there exists a list strong edge-coloring $\varphi^{\prime}$ of $G^{\prime}=G\setminus\left\{v_{1},v_{2},v_{3},v_{4},v_{5}\right\}$ for any list assignment $L$ with lists of size at least $10$. Let $\varphi$ be the coloring of $G$ induced by $\varphi^{\prime}$. Then, only the edges of $C$ and the pendant edges except $y_{0}$ are non-colored in $\varphi$. The edges $x_{0},x_{1},x_{4}$ and $x_{5}$ have at least $5$ available colors, the edges $x_{2}$ and $x_{3}$ have at least $6$, $y_{1}$ and $y_{5}$ have at least $3$, and $y_{2},y_{3},y_{4}$ have at least $4$ available colors.
@@
$P_{C_{k}}^{0}(x_{0},x_{1},y_{0},y_{1})$ $=\mathrm{coef}(P_{C_{k}}^{k-1};\ x_{k-1}^{4}y_{k-1}^{3})\cdot$
$\cdot(x_{0}-x_{1})(x_{0}-y_{0})(x_{0}-y_{1})(x_{1}-y_{0})(x_{1}-y_{1})(y_{0}-y_{1})\,,$
@@ 4 Graphs $G$ with $\chi_{s}^{\prime}(G)<\chi_{s,l}^{\prime}(G)$
In this section, we give a negative answer to Question 1.5 by proving Theorem 1.6. First, we recall the result about cubic graphs with strong chromatic index equal to $5$. It uses the notion of covering graphs defined as follows. A surjective graph homomorphism $f\colon\tilde{G}\to G$ is called a covering projection if for every vertex $\tilde{v}$ of $\tilde{G}$ the set of edges incident with $\tilde{v}$ is bijectively mapped onto the set of edges incident with $f(\tilde{v})$. The graph $G$ is usually referred to as the base graph and $\tilde{G}$ as a covering graph or a lift of $G$. A graph $\tilde{G}$ covers $G$ if there exists such a covering projection.
@@
Let $G^{\prime}$ be the graph obtained from $G$ by removing all the edges of $P^{\prime}$ except $uv_{1}$ and $vv_{2}$. Clearly, $G^{\prime}$ is the graph $R$ with one edge removed and replaced with two pendant edges, and thus it admits a strong $5$-edge-coloring $\varphi^{*}$ induced by the coloring $\varphi$ of $R$. Note that in $\varphi$, the edges $uv_{1}$ and $vv_{2}$ receive the same color (the color $\varphi(uv)$). Now, we show that in any strong $5$-edge-coloring of $G^{\prime}$ these two edges must be colored with the same color. Since the edges of $G^{\prime}$ in $L$ have the same lists of size $5$, this will imply that the two edges must receive the same color in any strong $L$-edge-coloring.
First, observe that in $\varphi^{*}$ the only common color the vertices $u$ and $v$ are incident with is color $a=\varphi(uv)$. Let $b\neq a$ be a color incident with $u$, and $c\neq a$ a color incident with $v$. Let $k$ be the number of edges of color $b$ in $\varphi^{*}$. Since every edge of color $b$ is adjacent to edges of all other four colors, the edges of $G^{\prime}$ colored with $b$ in $\varphi^{*}$ together with their adjacent edges cover all the edges of $G^{\prime}$ (every edge exactly once) except $vv_{2}$; we denote this (almost) covering $C_{b}$. Similarly, the edges of $G^{\prime}$ colored with $c$ in $\varphi^{*}$ together with their adjacent edges cover all the edges of $G^{\prime}$ (every edge exactly once) except $uv_{1}$; we denote this (almost) covering $C_{c}$.
@@
Similarly, we can show that around the vertices $u$ and $v$ in $G$, all five colors appear (i.e., the only common incident color is the color of the edges $uv_{1}$ and $vv_{2}$). Observe that the edges of $G^{\prime}$ colored with $a$ in $\varphi^{*}$ (except the edges $uv_{1}$ and $vv_{2}$) together with their adjacent edges cover all the edges of $G^{\prime}$ (every edge exactly once) except the edges incident with $u$ and $v$; we denote this covering $C_{a}$. Again, in every strong $5$-edge-coloring of $G^{\prime}$, on the edges of $C_{a}$ every color appears $k-1$ times, while in the whole graph every color appears on $k$ edges, except the color of $uv_{1}$ and $vv_{2}$, which appears $k+1$ times. This means, that $u$ and $v$ together are incident with edges of all five colors.
Now consider the coloring of the edges of $P^{\prime}$. Clearly, in any strong $L$-edge-coloring, all the five colors from $\left\{1,2,3,4,5\right\}$ appear on the inner edges of $P^{\prime}$. Similarly, since every spoke edge of $P^{\prime}$ sees $4$ distinct colors on the inner edges of $P^{\prime}$, every spoke edge can be colored with precisely one of the colors from $\left\{1,2,3,4\right\}$ or color $6$, except for the spoke edge that does not have color $5$ in its $2$-edge-neighborhood—that edge must be colored with $6$. Moreover, the edges $v_{1}w_{1}$ and $v_{2}w_{2}$ must be colored with the same color, so that we can combine the colorings of $G^{\prime}$ and $P^{\prime}$.
There are three non-isomorphic possibilities on which inner edge color $5$ appears. First, suppose that $u_{1}u_{4}$ is colored with $5$. Then, $u_{5}v_{5}$ must be colored with $6$ and therefore $v_{2}v_{3}$ is the only outer edge of $P^{\prime}$ which can be colored with $5$ or $6$. Therefore the colors $1$, $2$, and $3$ must be used on the remaining outer edges, and consequently $v_{1}w_{1}$ and $v_{2}w_{2}$ must also both be colored with either $5$ or $6$. This is not possible, since $v_{1}w_{1}$ has both colors in it¡s $2$-edge-neighborhood.
Second, suppose that $u_{3}u_{5}$ is colored with $5$. Then, $u_{4}v_{4}$ must be colored with $6$ and thus no outer edge of $P^{\prime}$ can have color $6$. Since every outer edge of $P^{\prime}$ also has color $5$ in the $2$-edge-neighborhood, it follows that the remaining four outer edges must be colored with colors $1$, $2$, and $3$. This means that $v_{1}v_{5}$ and $v_{2}v_{3}$ receive the same color, say $1$. But then, color $1$ cannot be incident with $u$ and $v$, and consequently, $u$ and $v$ together will not be incident with all five colors, which is not possible by the argument above.
So, we may assume that $u_{1}u_{3}$ is colored with $5$. Then, $u_{2}v_{2}$ must be colored with $6$, and $v_{1}w_{1}$ and $v_{2}w_{2}$ must both be colored with the same color as $u_{3}u_{5}$, which cannot be $4$—say it is $1$. Then, the outer edges of $P^{\prime}$ must be colored with colors from $\left\{2,3,5,6\right\}$. Since only $v_{4}v_{5}$ can be colored with $5$, it follows that $v_{1}v_{5}$ must have color $6$. Therefore, $u_{2}v_{2}$ and $v_{1}v_{5}$ both have color $6$, which means that, since $uv_{1}$ and $vv_{2}$ both have color $1$, some color, different $1$, must be incident with $u$ and $v$. As we showed above, this is not possible, and therefore a strong $L$-edge-coloring of $G$ does not exist. ∎
@@
Now, we consider several cases regarding the possible colorings of the maximum induced matchings. Note that throughout the process of coloring the edges, as soon as some color is picked for an edge $e\in M_{i}$, this color is removed from the lists of the edges which are in conflict with $e$, i.e., the edges of the maximum induced matchings different from $M_{i}$.
@@ 5 Graphs $G$ with $\chi_{n}^{\prime}(G)<\chi_{n,l}^{\prime}(G)$
In this section, we consider the results on list normal edge-coloring. As already mentioned, lists of size at least $10$ are always enough to find a normal list edge-coloring of a cubic graph. We do not know whether this bound is tight; currently, there are only examples of graphs with list normal chromatic index equal to $9$.
@@
There is an infinite family of cubic graphs with list normal chromatic index at least $9$.
@@
There are bridgeless cubic graphs with list normal chromatic index at least $8$.
@@
The graph in the proof of Theorem 5.2 has a $2$-edge-cut. So, again a question arises whether there are cubic graphs with high list normal chromatic index and high connectivity. We focused on cyclically $4$-edge-connected cubic graphs and surprisingly there are such graphs with list normal chromatic index at least $7$.
@@
There is an infinite family of cyclically $4$-edge-connected cubic graphs with list normal chromatic index at least $7$.
@@ 6 Conclusion
One of the main results of this paper is the tight upper bound of $10$ colors for the strong chromatic index of subcubic graphs. However, since this bound is only known to be attained by the Wagner graph and graph containing the $K_{3,3}$ with a subdivided edge as a subgraph, one may ask whether there are other examples of such graphs, perhaps with smaller chromatic index.
@@
As the second main result, we proved that the strong chromatic index and the list strong chromatic index differ for some graphs; we provided an infinite family of such graphs, but the family only contains graphs with the minimum possible value of the strong chromatic index, and it does not seem likely that for graphs with strong chromatic index closer to the general upper bound of $10$ colors, their list strong chromatic index will be different. Therefore, we propose a rather bold statement, which is in line with Conjecture 1.3.
@@
The first step towards proving this conjecture would be proving that the list strong chromatic index of any connected bridgeless subcubic graph, not isomorphic to the Wagner graph, is at most $9$. Or even more specifically, finding the exact upper bounds for the list strong chromatic indices of special graph families such as planar graphs and bipartite graphs would also give a relevant insight into the topic.
@@
In the case of list normal edge-coloring, we have the upper bound given by Theorem 1.4, but we do not have an example of a graph attaining the bound; in fact, we do not believe one exists.
@@
Conjecture 1.8 assumes only bridgeless cubic graphs. We showed in Theorem 5.2 that in the list version, there are bridgeless cubic graphs with list normal chromatic index at least $8$. In the proof of the theorem, we only provided one graph of order $12$. However, we are only aware of two other graphs with list normal chromatic index at least $8$; namely the Wagner graph and the graph obtained from $K_{3,3}$ in which one vertex is truncated (see Figure 10). We also remark here without a proof that with some additional effort, one can show that the list normal chromatic index of the Wagner graph is equal to $8$.
Figure 10: Two bridgeless cubic graphs with list strong chromatic index at least $8$. If the bold edges receive unique lists of seven colors and all the other edges the same lists of seven colors, then one can not realize a list normal edge-coloring.
Based on our results and additional computer tests on small graphs, we confidently propose also the following.
@@
    cyclically $k$-edge-connected, for some integer $k$?